2 masses of 10kg and 20kg respectively are connected by a mass less spring which are placed vertically . A force of 200N is acting on 20kg mass at the instant when the 10kg mass has an acceleration of 12 metre per second square towards 20kg mass the acceleration of 20kg mass is
1
Expert's answer
2020-04-06T09:09:00-0400
"a = F\/m"

"a_{total} = \sqrt {a_1+a_2}"

"a = \sqrt {12 + 200/20} = \sqrt {22} = 4.69 m/s^2 \approx 5m/s^2"